OpenStudy (anonymous):

Find the specified vector or scalar. u = -4i + 1j and v = 4i + 1j; Find ‖u+v‖. A. √34 B. 8 C. 2 D. 5

OpenStudy (helder_edwin):

\[\large u+v=(-4i+j)+(4i+j)=2j \] so \[\large \lVert u+v\rVert=\sqrt{2^2}=2 \]

OpenStudy (anonymous):

Thank you!